### v1

# Unsupervised Visual Chain-of-Thought Reasoning via Preference Optimization

Abstract
Chain-of-thought (CoT) reasoning greatly improves the interpretability and problem-solving abilities of multimodal large language models (MLLMs). However, existing approaches are focused on text CoT, limiting their ability to leverage visual cues. Visual CoT remains underexplored, and the only work [28] is based on supervised fine-tuning (SFT) that relies on extensive labeled bounding-box data and is hard to generalize to unseen cases. In this paper, we introduce Unsupervised Visual CoT (UV-CoT), a novel framework for image-level CoT reasoning via preference optimization. UV-CoT performs preference comparisons between model-generated bounding boxes (one is preferred and the other is dis-preferred), eliminating the need for bounding-box annotations. We get such preference data by introducing an automatic data generation pipeline. Given an image, our target MLLM (e.g., LLaVA-1.5-7B) generates seed bounding boxes using a template prompt and then answers the question using each bounded region as input. An evaluator MLLM (e.g., OmniLLM-12B) ranks the responses, and these rankings serve as supervision to train the target MLLM with UV-CoT by minimizing negative log-likelihood losses. By emulating human perception–identifying key regions and reasoning based on them–UV-CoT can improve visual comprehension, particularly in spatial reasoning tasks where textual descriptions alone fall short. Our experiments on six datasets demonstrate the superiority of UV-CoT, compared to the state-of-the-art textual and visual CoT methods. Our zero-shot testing on four unseen datasets shows the strong generalization of UV-CoT. The code is available in https://kesenzhao.github.io/my_project/projects/UV-CoT.html.
1 Introduction
With the recent advancements in multimodal large language models (MLLMs) [17, 18, 3, 39], many efforts have been made to incorporate text CoT reasoning [34, 13, 42, 8] to handle complex vision-language tasks [43, 37, 36]. However, the visual understanding ability of MLLM is limited by fixed-granularity image processing, i.e., the MLLM can’t dynamically adjust focus across different spatial regions of the input image, even when guided by text CoT prompts [9]. This underscores the critical need to explicitly integrate visual cues into the CoT process.
Figure 1: Comparison of Visual CoT [28] and our UV-CoT. Left: Visual CoT relies on human-annotated bounding boxes to identify key regions. The model is trained via supervised fine-tuning to maximize the likelihood of the labeled data. Right: UV-CoT eliminates the need for human annotation. Given an image, the target model generates seed bounding boxes and answers questions based on these regions, respectively. An evaluator MLLM then scores the responses as a proxy for assessing region quality. Lastly, the target model is optimized via preference optimization by maximizing the likelihood of preferred regions over dispreferred ones.
A very recent work called Visual-CoT [28] has made an initial attempt towards this goal. The model is trained using supervised fine-tuning (SFT) with human-labeled bounding boxes that indicate the key image regions relevant to the question. Following this, it performs its multimodal CoT approach with human-annotated reasoning steps by associating textual inputs with the detected regions. An overview of Visual CoT is presented in Fig. 1. However, it has two key drawbacks: (1) it relies on large-scale, high-quality labeled data, making it costly and hard to scale; and (2) SFT learns only from positive examples (i.e., the labeled training data), limiting its ability to generalize to unseen or ambiguous scenarios where intermediate reasoning or dynamic interpretation is needed.
To address these issues, we introduce an Unsupervised approach to Visual CoT dubbed as UV-CoT. It has two key parts: data collection and model training. The data collection does not need human annotation, as it leverages the data generation and evaluation capabilities of pre-trained MLLMs. The model training is inspired by the idea of direct preference optimization (DPO). It is implemented with an adapted version of DPO to address specific limitations in capturing the degree of preference and fine-grained region-based reasoning when conducting visual CoT on MLLMs.
Our method UV-CoT, as shown in Fig. 1, differs from Visual CoT [28] by adopting an unsupervised approach with contrastive preference data. We design an automatic two-step pipeline to generate this data: 1) Region Generation: Given an image, the target model generates multiple seed bounding boxes using a template prompt. It then answers the question by processing each bounded region together with the question as input. 2) Quality Assessment: An evaluator MLLM scores the responses, using these scores as proxies to measure the quality of the regions. Unlike traditional DPO, we propose Score-DPO (sDPO), which not only ranks preference data (i.e., preferred and dis-preferred responses shown in Fig. 1) but also assigns preference scores. This scoring enables more precise optimization based on score differences. During UV-CoT training, the rankings of the preference data act as supervision by minimizing negative log-likelihood loss, while the scores define the decision margin. By mimicking human perception–first identifying key regions, then reasoning over them–UV-CoT significantly improves visual comprehension, especially in spatial reasoning tasks where text-based methods fall short. By leveraging unsupervised data in a contrastive way, UV-CoT also shows strong generalization ability when tested on unseen datasets. We will show the evidence in the experimental sections.
Our contributions in this paper thus include: 1) an automatic pipeline for generating high-quality preference data, enabling robust and scalable preference learning of UV-CoT; 2) an improved version of DPO in UV-CoT by integrating the degree of preference for visual regions, allowing the model to distinguish key regions more precisely and improve its reasoning ability; and 3) extensive experiments on multiple challenging datasets, demonstrating state-of-the-art performance of UV-CoT on four benchmarks and strong generalization to four unseen test datasets.
2 Related Works
Chain-of-Thought in LLMs and MLLMs. LLMs [19, 32, 3] show strong inferential abilities, with CoT [34, 7] improving their performance by guiding models through intermediate reasoning steps. Both manually designed [34] and self-generated [13, 42] reasoning approaches have proven effective in improving LLM performance. In contrast, MLLMs struggle with reasoning due to differences in processing textual and visual data. Multimodal CoT methods [43, 37, 36] attempt to address this by transforming multimodal inputs into a unified textual format, enabling LLMs to perform CoT at the text level. However, this transformation introduces significant information loss and prevents the models from capturing key visual details [43]. For example, LLaVA-CoT [36] leverages GPT-4o to summarize questions and image captions but suffers from weak optical character recognition and sometimes hallucinations.
A very recent work, Visual CoT [28], improves the MLLM reasoning by introducing CoT methods at the image level. This approach involves scanning the entire image, identifying key references, and then focusing the model on specific regions for reasoning. Despite its improvements, Visual CoT is heavily based on costly human-annotated data. In contrast, our UV-CoT framework utilizes unsupervised preference optimization with auto-generated preference data, eliminating the need for manual annotations.
Preference Learning in LLMs and MLLMs. RLHF [31] aligns LLMs with human preferences by training a reward model via contrastive response evaluations [45]. To reduce reliance on human annotations, RLAIF [15] leverages pre-trained LLMs for preference label generation. However, RL-based fine-tuning faces stability and efficiency challenges. Direct Preference Optimization (DPO) [26] addresses this by directly linking reward functions to optimal policies, eliminating reward model fine-tuning. Further improvements include IPO [2], which mitigates overfitting with a bounded preference function, and KTO [6], which removes the need for paired preference data, relying instead on single examples labeled as either “good” or “bad”.
These preference learning techniques are applied to MLLMs, with RLHF-V and RLAIF-V [40, 41] refining behavior alignments using human and LLM-generated labels. To mitigate reward hacking, Sun et al. [30] enhance reward models with additional factual details, such as image captions and verified choices, further improving the performance of MLLM. Some works have attempted to apply DPO in the CoT process [24, 14]. However, these methods are designed for only text-level CoT and do not effectively handle visual features or cues. In contrast, in this paper, we propose UV-CoT, a specialized framework for image-level CoT reasoning inspired by the idea of DPO. Unlike traditional DPO, UV-CoT not only ranks preference data (i.e., preferred and dis-preferred data) but also assigns preference scores. This scoring enables more precise optimization of the MLLMs based on score differences.
3 Method
Figure 2: Illustration of UV-CoT reasoning.
Fig. 1 illustrates the pipeline of UV-CoT reasoning. Given the original image and the question, we append a CoT prompt to guide the target MLLM in identifying the most informative image region and specifying its location via bounding box coordinates. A visual sampler then extracts the bounded region from the image. The MLLM subsequently integrates visual tokens from original and cropped images to generate more precise and comprehensive answers. In Sec. 3.1, we detail the automatic preference data generation pipeline. In Sec. 3.2, we describe our Score-DPO (sDPO) for image-level CoT reasoning.
Algorithm 1 Preference data generation for a query $x$
1:Input: Target model $f_{\mathsf{tar}}$, evaluator $f_{\mathsf{eval}}$, an image-question query $x$, number of seeds $n$, and number of preference pairs $k$.
2:Output: Preference data $\mathcal{D}$
3:Initialize $y_{0}=x$
4:for $t=1$ to $T$ do
5:     $\{y_{t}^{i}\}_{i=1}^{n}\leftarrow\mathsf{Generate}(f_{\mathsf{tar}},y_{0:t-1},n)$
6:     $\{s^{i}\}_{i=1}^{n}\leftarrow\mathsf{Evaluate}(f_{\mathsf{eval}},y_{0:t-1},\{y_{t}^{i}\}_{i=1}^{n})$
7:     $\mathcal{D}_{t}\leftarrow\mathsf{ConstructPairs}(y_{0:t-1},\{y_{t}^{i}\}_{i=1}^{n},\{s^{i}\}_{i=1}^{n},k)$
8:     $y_{t}\leftarrow\mathsf{Select}(y_{0:t-1},\{y_{t}^{i}\}_{i=1}^{n},\{s^{i}\}_{i=1}^{n})$
9:end for
10:return $\mathcal{D}=\{\mathcal{D}_{1},\dots,\mathcal{D}_{T}\}$
3.1 Preference Data Generation
Given a target model $f_{\mathsf{tar}}$, an evaluator model $f_{\mathsf{eval}}$, and an image-question pair $x$, we illustrate how to construct $n$ preference data points. The evaluator model $f_{\mathsf{eval}}$ is a slightly stronger MLLM than the target model $f_{\mathsf{tar}}$, providing a better assessment of the responses [^1] Please note that the target MLLM can also be used as the evaluator, validated in Tab. 5.. Assume we have $T$ steps in CoT reasoning, we generate preference data for $T$ times on the way, as described in Algorithm 1. At each timestep $t$ (i.e., a reasoning step $t$), the process includes four stages: Response Generation, Response Evaluation, Pair Construction, and Response Selection.
Response Generation. The goal of this stage is to generate seed bounding boxes and produce intermediate responses to the question using the target model. Here, a “response” refers to any model output at an intermediate step, not necessarily the final answer to the question. We denote the model’s response at timestep $t$ as $y_{t}$, with the initial input $x$ represented as $y_{0}$. To create diverse bounding boxes and subsequent responses, we apply stochastic decoding on the target model $f_{\mathsf{tar}}$ with $n$ different seeds, resulting in a set of responses ${y_{t}^{i}}_{i=1}^{n}$.
Response Evaluation. This stage evaluates the quality of all generated responses. The evaluator model not only scores each individual response but also counts how this response influences the quality of its next response in the chain. This cumulative evaluation approach helps quantify the impact of each bounded region on the overall reasoning process. Below, we elaborate on the formulation. At timestep $t$, the evaluator assign scores for $y_{t}^{i}$ as follows:
$s_{\mathsf{cur}}^{i}$ $=f_{\mathsf{eval}}(y^{i}_{t}\mid y_{0:t-1}),$ (1)
$s_{\mathsf{nxt}}^{i}$ $=\mathbb{E}[f_{\mathsf{eval}}(y_{t+1}^{1}\mid y_{0:t-1},y^{i}_{t})],$
$s^{i}$ $=s_{\mathsf{cur}}^{i}+\gamma s_{\mathsf{nxt}}^{i},$
where $s_{\mathsf{nxt}}^{i}$ reflects the impact on the next response and $\gamma>0$ is a hyperparameter to combine the scores of current and the next response, with $\gamma=0$ at the last step. We estimate the expectation $\mathbb{E}[\cdot]$ by randomly sampling next responses.
Pair Construction. At each timestep $t$, we randomly select $k$ (preferred and dis-preferred) pairs from $\{y_{t}^{i}\}_{i=1}^{n}$. For a single pair, we concatenate the preferred response with the past response chain $y_{0:t-1}$ (preserved after $t\!-\!1$ timesteps) and get a “preferred chain” denoted as $y_{t}^{w}$, and then we concatenate the dis-preferred response in the same way to get a “dis-preferred chain” denoted as $y_{t}^{l}$. The pair of chains also includes their respective scores, and they are represented as $\{y_{w},s_{w},y_{l},s_{l}\}$. The overall $k$ pairs of chains compose the preference dataset $\mathcal{D}_{t}$.
Response Selection. The abovementioned “past response chain $y_{0:t-1}$” is unique and is concatenated by the highest-scoring response at timestep $t\!-\!1$ and the preserved chain at timestep $t\!=\!2$, i.e., $y_{0:t-2}$. In other words, when finishing each timestep process, we preserve only the best chain and use it for the next step.
3.2 Unsupervised Learning of UV-CoT
Assume the preference dataset $\mathcal{D}$ has been generated across $t$ timesteps, we then optimize the target model with our UV-CoT via preference optimization on $\mathcal{D}$. DPO [1] is widely used in preference learning, and it ranks responses without quantifying preference intensity. In our case of image-level reasoning, key regions vary in influence, necessitating finer differentiation between responses. Therefore, we refine DPO by adjusting the margin to capture the key region’s influence. We name our loss Score-DPO, abbreviated as sDPO, as it incorporates the preference score into the optimization. The loss is formulated as:
$\mathcal{L}_{sDPO}(\theta)=-\underset{\left(x,y_{w},y_{l}\right)\sim\mathcal{D}}{\mathbb{E}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}\mid x)}{\pi_{\mathrm{ref}}(y_{w}\mid x)}\right.\right.$ (2)
$\quad\left.\left.-\beta\log\frac{\pi_{\theta}(y_{l}\mid x)}{\pi_{\mathrm{ref}}(y_{l}\mid x)}-(g(s_{w})-g(s_{l}))\right)\right],$
where $\pi_{\theta}$ is the target model, and $\pi_{\text{ref}}$ is its frozen initialization, serving as a reference to constrain deviation from the original model. $g(\cdot):\mathbb{R}\rightarrow\mathbb{R}$ is a monotonically increasing function that maps preference scores into the logit space of the DPO objective.
To provide a deeper understanding of our sDPO loss, we establish its connection to the standard DPO loss. DPO reformulates reward model training as policy optimization by reparameterizing the reward function in PPO [27]:
$r(x,y)=\beta\log\frac{\pi_{\theta}(y\mid x)}{\pi_{\mathrm{ref}}(y\mid x)}+\beta\log Z(x)$ (3)
Using the Bradley-Terry [4] model, the standard DPO aims to minimize the negative log-likelihood of the difference of rewards:
${P}\left(y_{w}-y_{l}>0\right)=\sigma\left(r\left(x,y_{w}\right)-r\left(x,y_{l}\right)\right)$ (4)
$\mathcal{L}_{\text{DPO}}=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log{P}\left(y_{w}-y_{l}>0\right)\right],$
where $\sigma(x)=\frac{1}{1+\exp(-x)}$ is the sigmoid function.
Let $\Delta_{r}=g(s_{w})-g(s_{l})$ and define the Gumbel-distributed random variables $R_{w}\sim\operatorname{Gumbel}\left(r\left(x,y_{w}\right),1\right)$ and $R_{l}\sim\operatorname{Gumbel}\left(r\left(x,y_{l}\right),1\right)$. Then, we derive:
$P\left(R_{w}-R_{l}>\Delta_{r}\right)=\sigma\left(r\left(x,y_{w}\right)-r\left(x,y_{l}\right)-\Delta{r}\right)$ (5)
$=\sigma\left(\beta\log\frac{\pi^{*}\left(y_{w}\mid x\right)}{\pi_{\mathrm{ref}}\left(y_{w}\mid x\right)}-\beta\log\frac{\pi^{*}\left(y_{l}\mid x\right)}{\pi_{\mathrm{ref}}\left(y_{l}\mid x\right)}-\Delta{r}\right)$
This result follows from the definition of Gumbel random variables [1] and the Gumbel-max trick [20]. We provide a detailed proof in the Appendix B. By maximizing the log-likelihood of this objective, we obtain our proposed loss function. The Gumbel distribution models the extreme values of a variable, while $\Delta_{r}$ quantifies the degree of difference between preference pairs. Thus, our loss function explicitly optimizes preference learning by distinguishing not only the order but also the magnitude of preference differences.
Algorithm 2 Iterative learning of UV-CoT
1:Input: Initial target model $f^{1}_{\mathsf{tar}}$, evaluator model $f_{\mathsf{eval}}$, $\mathcal{X}=\{\mathcal{X}_{1},\dots,\mathcal{X}_{m}\}$, each $\mathcal{X}_{i}$ is a subset of image-question query
2:Output: $f^{m}_{\mathsf{tar}}$
3:for $i=1$ to $m$ do
4:     $\mathcal{D}_{i}\leftarrow\mathsf{GenerateData}(f^{i}_{\mathsf{tar}},f_{\mathsf{eval}},\mathcal{X}_{i})$triangleright Algorithm 1
5:     $f^{i+1}_{\mathsf{tar}}\leftarrow\mathsf{Train}(f^{i}_{\mathsf{tar}},\mathcal{D}_{i})$triangleright  Eq. 2
6:end for
7:return $f^{m}_{\mathsf{tar}}$
Iterative Learning. Standard DPO relies on static preference data during training, which can lead to distributional mismatch between training data and the model’s generated outputs. To mitigate this issue, we adopt an iterative learning approach inspired by [41]. Algorithm 2 outlines the iterative learning process of UV-CoT, which incrementally refines a target model through preference learning. The iteration repeats $m$ times, and the total image-question query set $\mathcal{X}$ is evenly divided into $m$ subsets, $\mathcal{X}=\{\mathcal{X}_{1},\ldots,\mathcal{X}_{m}\}$ , each assigned to one iteration. The process starts with an initial target model $f_{\mathsf{tar}}$, an evaluator model $f_{\mathsf{eval}}$, and $\mathcal{X}$. In each iteration $i$, the algorithm first generates preference data $\mathcal{D}_{i}$ using the current target model $f^{i}_{\mathsf{tar}}$ and the subset $\mathcal{X}_{i}$, using the procedure in Algorithm 1. This newly generated preference data is then used to train the next iteration of the target model $f^{i+1}_{\mathsf{tar}}$ using our UV-CoT loss of Eq. 2. The process continues until the final model $f^{m}_{\mathsf{tar}}$ is obtained. By dynamically updating the preference data, our approach ensures that the learned model adapts to its evolving distribution, enhancing training robustness.
MLLM DocVQA TextVQA InfographicsVQA Flickr30k GQA VSR Average LLaVA-1.5-7B 0.198 0.507 0.131 0.539 0.480 0.504 0.393 LLaVA-1.5-13B 0.225 0.543 0.169 0.607 0.506 0.512 0.427 MiniCPM-o-8B 0.232 0.529 0.175 0.618 0.495 0.521 0.428 OmniLMM-12B 0.254 0.578 0.172 0.621 0.509 0.523 0.443 Visual-CoT-7B (100% label) 0.294 0.673 0.194 0.652 0.546 0.532 0.482 UV-CoT (0% label) 0.265 0.686 0.173 0.632 0.536 0.548 0.473 UV-CoT (10% label) 0.283 0.711 0.198 0.649 0.568 0.553 0.494
Table 1: Overall comparison of different models on six evaluation benchmarks. The best result is bold, the second-best is underlined. “%” indicates the percentage of supervised data used in Visual-CoT. By default, our UV-CoT uses only unsupervised data.
4 Experiments
MLLM DUDE SROIE Visual7w Average LLaVA-1.5-7B 0.165 0.147 0.340 0.217 LLaVA-1.5-13B 0.174 0.159 0.352 0.228 MiniCPM-o-8B 0.182 0.165 0.341 0.229 OmniLMM-12B 0.194 0.166 0.357 0.239 Visual-CoT-7B 0.206 0.181 0.397 0.261 UV-CoT 0.241 0.184 0.432 0.286 UV-CoT 0.253 0.227 0.455 0.312
Table 2: Zero-short experiments on DUDE, SROIE and Visual7w. The best result is bold, the second-best is underlined. ‘UV-CoT’ denotes our model trained with additional unlabeled preference data from the three zero-shot datasets.
MLLM Attributes GPT4V-hard OCR Average LLaVA-1.5-7B 0.317 0 0.1 0.139 LLaVA-1.5-13B 0.326 0.118 0.133 0.192 MiniCPM-o-8B 0.322 0.118 0.1 0.180 OmniLMM-12B 0.326 0.118 0.167 0.204 Visual-CoT-7B 0.330 0.118 0.593 0.347 UV-CoT 0.352 0.176 0.677 0.402
Table 3: High-resolution image reasoning on V Bench. The best result is bold.
Model DocVQA TextVQA InfographicsVQA Flickr30k GQA VSR Average UV-CoT (10% labels) 0.283 0.711 0.198 0.649 0.568 0.553 0.494 IF: w/o UV-CoT 0.149 0.574 0.160 0.585 0.509 0.522 0.417 IF: UV-CoT w/ G.T. BBox 0.528 0.769 0.504 0.655 0.664 0.585 0.618 Tr: w/ naive DPO 0.258 0.695 0.189 0.623 0.552 0.534 0.475 Tr: w/o iterative learning 0.256 0.671 0.162 0.609 0.523 0.531 0.459 Ge: w/o $\gamma$ 0.247 0.539 0.152 0.551 0.484 0.460 0.406
Table 4: Ablation study on key components of UV-CoT (10% labeled data). ‘w/o UV-CoT’ denotes standard inference without CoT reasoning. ‘UV-CoT w/ G.T. BBox’ uses annotated ground truth bounding boxes. ‘w/ naive DPO’ applies the standard DPO loss. ‘w/o iterative learning’ generates preference pairs for the entire set of question queries $\mathcal{X}$ in a single pass and trains once. ‘w/o $\gamma$’ evaluates responses with $\gamma=0$. IF, Tr, and Ge indicate ablations on the inference process, training loss, and data generation, respectively.
In Sec. 4.1, we outline our experimental setup, including datasets, evaluation metrics, baselines, and implementation details. Sec. 4.2 evaluates UV-CoT on six widely used datasets, demonstrating its superiority over existing baselines. Sec. 4.4 presents ablation studies analyzing impact of key components on overall performance. In Sec. 4.3, we assess the generalization ability of UV-CoT in zero-shot settings across four diverse datasets. In Sec. 4.5, we evaluate the quality of generated bounding boxes across all datasets. In Sec. 4.6, we analyze how our evaluation process reduces task difficulty. In Sec. 4.7, we provide additional analyses on token efficiency, visualization.
4.1 Experimental Setup
Datasets. To generate diverse image-level CoT data, we select nine datasets spanning five domains:
  • Text/Document: DocVQA [21], TextVQA [29], DUDE [33], and SROIE [10].
  • Chart: InfographicsVQA [22].
  • General VQA: Flickr30k [25] and Visual7W [44].
  • Relation Reasoning: VSR [16] and GQA [11].
  • High-Resolution Image Reasoning: V Bench [35].
We follow the procedure in Sec. 3.1 to generate high-quality preference data based on these datasets. Notably, we also provide a model trained on data excluding DUDE, SROIE, and Visual7W, which is used to evaluate zero-shot performance. For details on dataset composition and preprocessing, please refer to Appendix C.1.
Evaluation. Following the evaluation methodology of prior work [18], we use GPT-4o [23] as the evaluator. GPT-4o is prompted to assign a numerical score between 0 and 1, with higher scores indicating better prediction accuracy. Details on prompt design and setup are provided in Appendix C.2.
Baselines We compare our model against five state-of-the-art baselines. Specifically, LLaVA-1.5-7B [17], LLaVA-1.5-13B and OmniLMM-12B serve as strong comprehensive baselines, as they are pretrained on large-scale multimodal data and fine-tuned with high-quality instruction data. OmniLMM-12B is also our evaluator model. Additionally, MiniCPM-o-8B [38] employs an adaptive visual encoding mechanism to capture fine-grained features, improving visual understanding. Finally, Visual-CoT-7B [28] utilizes labeled data to learn image-level CoT via SFT, achieving competitive performance.
Implementation Details. We use LLaVA-1.5-7B as the target model and OmniLMM-12B as the evaluator. To ensure scalability, we avoid using GPT-4 as the evaluator, preventing constraints imposed by high API costs. Our UV-CoT training follows a two-stage process aligned with the LLaVA-1.5 training framework. In the first stage, the vision encoder and language model weights are frozen while training on image-text caption data. In the second stage, all weights are unfrozen for full training. We implement iterative learning of UV-CoT over four iterations, utilizing a total of 249K preference data pairs. Notably, UV-CoT achieves higher data efficiency than Visual-CoT, which uses 376K data pairs. For each iteration, we train the model with AdamW optimizer for 4 epochs with a learning rate of $0.15\times 10^{-7},\beta=$ 0.1, and a batch size of 8. In total, data generation takes 80 hours and training requires 60 hours, both conducted on an 8$\times$×A100 40GB machine. Additionally, we provide a variant UV-CoT trained with extra SFT on 10% of the labeled Visual-CoT data, denoted as UV-CoT (10%).
4.2 Comparisons to State-of-The-Arts
The overall performance comparisons are reported in Tab. 1, leading to the following key observations:
  • Both image-level CoT models, Visual-CoT and our UV-CoT, achieve a significant performance gain over all baselines, even outperforming the larger LLaVA-1.5-13B model across all datasets. This highlights the effectiveness of image-level CoT reasoning. Although MiniCPM-o-8B employs crop-and-select strategies to enhance focus on specific image regions, this approach remains limited. In contrast, image-level CoT fully utilizes MLLMs to generate key regions, whereas MiniCPM-o-8B relies on rule-based cropping applied to only a few layers. Nonetheless, MiniCPM-o-8B outperforms LLaVA-1.5-7B by an average of 3.5%, underscoring the importance of explicitly integrating visual cues.
  • UV-CoT significantly outperforms OmniLMM-12B by an average of 5.1% across all evaluated datasets. This underscores that our method extend beyond merely distilling knowledge from larger evaluator model. Directly generating accurate bounding boxes is a challenging task for MLLMs, requiring precise spatial understanding and localization. UV-CoT addresses this challenge by reformulating the task to focus on evaluating the final answer, which is inherently more tractable.
  • UV-CoT outperforms Visual-CoT-7B on TextVQA (+1.3%) and VSR (+1.6%). Moreover, UV-CoT (10%) surpasses Visual-CoT-7B across four datasets and 2.1% in average, with particularly notable improvements on TextVQA (+2.5%), GQA (+2.2%), and VSR (+2.1%). On the remaining two datasets, UV-CoT (10%) achieves comparable performance to Visual-CoT-7B. Notably, our method is trained on significantly less unlabeled data (249K) than Visual-CoT-7B, which uses 376K labeled CoT data. This success can be attributed to our high-quality preference data generation and the enhanced preference optimization method, which further improves image-level CoT learning.
4.3 Zero-Shot Performance (Generalization)
Comprehensive Zero-Shot. We evaluate our model’s zero-shot image-level CoT capabilities using the test splits of SROIE, DUDE, and Visual7W, without training on these datasets. Since our method does not require labeled data, we can effectively evaluate these datasets even in the absence of bounding box annotations. Specifically, we train a variant of our model, denoted as UV-CoT, using preference data that incorporates these datasets. The results are presented in Tab. 2. UV-CoT consistently outperforms all baselines across all zero-shot datasets (2.5% in average), achieving notable improvements on DUDE (+3.5%) and Visual7W (+3.5%), which underscores its strong generalization ability. In contrast, supervised CoT learning, which depends on labeled data, often overfits to specific annotation distributions, limiting its ability to generalize to new tasks. UV-CoT, however, leverages preference optimization by focusing on relative comparisons rather than absolute labels, thereby reducing reliance on specific annotated distributions. Furthermore, UV-CoT achieves even greater improvements (5.1% in average), with a particularly significant margin on SROIE (+4.6%), demonstrating that our model can effectively learn the image-level CoT process without depending on labeled data. Based on this capability, our method can be easily extended to additional tasks, eliminating the need for costly human annotation.
High-Resolution Image Reasoning. To evaluate the capability in handling challenging high-resolution image reasoning tasks, we test UV-CoT on the V Bench [35], which comprises 238 images from the SA-1B dataset [12] with an average resolution of 2246$\times$1582. We conduct experiments on three representative tasks—Attributes (object attribute recognition), GPT4V-Hard (complex visual reasoning), and OCR (text recognition)-with the results summarized in Tab. 3. CoT-based methods, UV-CoT (+19.8% on average) and Visual-CoT-7B (+14.3% on average), significantly outperform non-CoT baselines, especially for OCR tasks (more than 50%). And UV-CoT achieves the highest performance across all tasks, surpassing Visual-CoT-7B by 5.5% on average. This illustrates the effectiveness of the visual CoT in high-resolution image reasoning. The superior performance over Visual-CoT-7B further demonstrates that UV-CoT optimizes the visual CoT learning process, enabling fine-grained region-based reasoning.
4.4 Ablation Studies
This section presents ablation experiments analyzing key components, as shown in Tab. 4.
Image-Level CoT: To evaluate the effectiveness of image-level CoT reasoning, we introduce two variants. w/o UV-CoT removes the CoT reasoning process, relying on direct inference without intermediate reasoning steps. UV-CoT w/ G.T. BBox utilizes ground-truth (GT) bounding boxes in the first step of the CoT process to assess the impact of accurate region selection. Results show that w/o UV-CoT leads to a significant performance drop across all datasets (7.7% in average), demonstrating the effectiveness of our image-level CoT approach. Meanwhile, UV-CoT w/ G.T. BBox serves as an upper bound for our method. In Flickr30k, our model performs on par with UV-CoT w/ G.T. BBox, indicating that it accurately selects key regions. However, in DocVQA and Infographics VQA, UV-CoT w/ G.T. BBox significantly outperforms our method, highlighting the challenge of generating precise bounding boxes. This suggests that our approach has a high upper limit and strong potential. Future work can focus on improving bounding box accuracy, especially for complex tasks.
Preference Optimization: w/ naive DPO applies a standard DPO loss. Its performance deteriorates across all datasets (1.9% in average), with particularly significant drops on DocVQA (2.5%) and Flickr30k (2.6%), highlighting the limitations of standard DPO in image-level CoT learning. This result highlights the advantages of sDPO, which enhances preference differentiation by incorporating scores. The consistent performance gains further validate the accuracy of our preference scores.
Iterative Learning: w/o iterative learning generates preference pairs for the entire set of question queries $\mathcal{X}$ in a single pass and trains once, leading to a significant performance drop (3.5% in average) across evaluated datasets. This underscores the importance of iterative learning in dynamically adapting the preference data distribution to remain aligned with the model’s evolving policy throughout the training process.
Response Evaluation: We assess the impact of the response evaluation stage by setting $\gamma=0$, resulting in the w/o $\gamma$ model. In this setting, the generated preference scores ignore the influence of the next response. This leads to a significant performance drop across all datasets (8.8% in average), particularly on TextVQA (17.2%), highlighting the challenge that MLLMs face in directly evaluating bounding boxes. This result underscores the necessity of incorporating next response in our evaluation method.
Figure 3: Object detection performance comparisons on (a) training datasets and (b) zero-shot datasets.
Model DocVQA TextVQA InfographicsVQA Flickr30k GQA VSR Average OmniLMM-12B 0.254 0.578 0.172 0.621 0.509 0.523 0.443 IF: OmniLMM-12B + CoT 0.305 0.722 0.217 0.675 0.580 0.595 0.516 LLaVA-1.5-7B 0.198 0.507 0.131 0.539 0.480 0.504 0.393 IF: LLaVA-1.5-7B + CoT 0.245 0.577 0.149 0.606 0.530 0.533 0.440 Tr: UV-CoT (Evaluator: self-evaluated) 0.242 0.609 0.153 0.598 0.517 0.526 0.441 Tr: UV-CoT (Evaluator: OmniLMM-12B) 0.265 0.686 0.173 0.632 0.536 0.548 0.473
Table 5: Analysis of evaluator model. + CoT refers to inference assisted by bounding boxes generated by UV-CoT. Self-evaluated denotes using target model as the evaluator during training. IF and Tr indicate experiments conducted on the inference and training process, respectively.
Figure 4: The model performance comparison regarding different sizes of visual tokens.
Figure 5: Visualization of preference data generated by Algorithm 1. Preferred bounding boxes are shown in red. Dis-preferred bounding boxes are in blue.
Figure 6: Visualization of our UV-CoT inference. Model-generated bounding boxes are shown in red.
4.5 Object Detection Performance
Previous experiments have demonstrated the exceptional performance of our image-level CoT process, highlighting the effectiveness of the generated bounding boxes in improving subsequent answers. However, we remain interested in evaluating the quality of the bounding boxes themselves. To this end, we employ GPT-4o to score the model-generated bounding boxes on both training datasets (Fig. 3 (a))and zero-shot datasets (Fig. 3 (b)), where best performances are marked with stars. InforVQA is a short for InfographicsVQA. We have the following observations:
  • Our UV-CoT outperforms Visual-CoT-7B, achieving higher scores in five out of six training datasets, supporting its superior performance in generating helpful bounding box.
  • The bounding box quality is closely related to the final performance. Our model exhibits lower scores (below 0.210) for bounding box generation in DocVQA and InfographicsVQA, correlating with its reduced final scores in these datasets ((below 0.290 in Tab. 1). It underscores the validity of evaluating bounding box quality through its impact on subsequent answers. In addition, this observation explains the performance gap between UV-CoT and UV-CoT w/ G.T. BBoxes in DocVQA and InfographicsVQA, illustrating the upper of our methods.
  • Both UV-CoT and UV-CoT perform better than Visual-CoT-7B across all zero-shot datasets. This demonstrates the strong generalization ability of our method in bounding box generation.
4.6 Insight of Evaluator Model
To gain a deeper understanding of the evaluator model, we conduct a series of variant experiments. The key findings, as presented in Tab. 5, are summarized as follows:
  • We compare UV-CoT with its self-evaluated variant, where the evaluator model is the same as the target model (initialed with LLaVA-1.5-7B). Although the self-evaluated version exhibits a performance decrease of 3.2% compared to the original UV-CoT, it still outperforms LLaVA-1.5-7B (+4.8% on average) across all evaluated datasets and achieves performance comparable to the larger OmniLMM-12B model (-0.2% on average). This demonstrates that UV-CoT maintains robust performance even under self-evaluation, highlighting its efficiency without requiring larger model scales.
  • For OmniLMM-12B and LLaVA-1.5-7B, we incorporate a CoT process using bounding boxes generated by UV-CoT. The CoT-enhanced versions significantly outperform their original counterparts, achieving average performance gains of 4.7% for LLaVA-1.5-7B and 7.4% for OmniLMM-12B across datasets. Remarkably, these models were not fine-tuned for the CoT process, indicating that the bounding box information alone substantially improves performance. This finding underscores that our evaluating process simplifies the task of generating complex spatial annotations, enabling MLLMs to focus on evaluating final answers.
4.7 Other Detailed Analyses
Visual Token Efficiency. Compared to standard answer generation in MLLMs, Visual CoT [28] and our UV-CoT reasoning double the number of visual tokens by processing additional local image regions. To evaluate token efficiency, i.e., performance under the same visual token budget, we resize the input image to different resolutions ($224^{2},336^{2}$ and $448^{2}$) and report the average score of different models in Fig. 4. (1) MLLMs with image-level reasoning (Visual CoT and our UV-CoT) demonstrate better token efficiency than the standard answer generation pipeline. Specifically, they achieve higher performance with 512 visual tokens than the standard pipeline does with 1024 tokens. (2) Our UV-CoT consistently outperforms the Visual CoT pipeline across all scales, achieving higher average scores due to it well-designed unsupervised learning pipeline. This highlights the token efficiency of our method.
Visualization. This section visualizes preference data and selected cases from our UV-CoT inference. As shown in Fig. 5, given different local regions and their corresponding answers, the evaluator MLLM assigns reasonable scores, validating the effectiveness of our automatic generation and labeling process. In Fig. 6, we present selected reasoning cases with model-generated bounding boxes overlaid. The precision of bounding box detection and the depth of understanding play a crucial role in determining the quality of generated answers.
5 Conclusion
In this work, we propose UV-CoT, a novel framework that enables image-level CoT reasoning in MLLMs via preference optimization techniques. Unlike previous methods that rely on SFT needing large amounts of labeled data, our approach leverages an unsupervised learning method to refine the model’s ability with image-level CoT using model-generated preference data (which are rough but useful). We address key challenges in preference data generation and effective optimization, ensuring a more adaptive and interpretable reasoning process. Extensive experiments demonstrate that UV-CoT achieves state-of-the-art performance, significantly improving visual comprehension in MLLMs on four reasoning datasets. Our findings from extensive experiments highlight the potential of preference learning as a scalable alternative to traditional SFT, enabling more robust and data-efficient multimodal reasoning.
Appendix A Outline
We begin by presenting an overview of our Appendix.
  • Section B: Framework Datails. We provide a detailed explanation on the connection between our UV-CoT and DPO.
  • Section C: Implement Details. We describe the specifics of our dataset, and evaluation methodology.
  • Section D: Limitations. We analysis the constraints and challenges of our approach.
  • Section E: Potential negative societal impacts. We discuss possible negative consequences and ethical considerations associated with our work.
Appendix B Connection between UV-CoT and DPO
B.1 Loss Function Formulation
To better captures the impact of key regions, we introduce a preference-weighted optimization approach. The loss function for UV-CoT is defined as follows:
$\mathcal{L}_{{sDPO}}(\theta)$ $=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\Bigg{[}\log\sigma\Bigg{(}\beta\log\frac{\pi_{\theta}(y_{w}\mid x)}{\pi_{\mathrm{ref}}(y_{w}\mid x)}$ (6)
$\quad-\beta\log\frac{\pi_{\theta}(y_{l}\mid x)}{\pi_{\mathrm{ref}}(y_{l}\mid x)}-\big{(}g(s_{w})-g(s_{l})\big{)}\Bigg{)}\Bigg{]}.$
where $\pi_{\theta}$ is the target policy model being optimized, $\pi_{\mathrm{ref}}$ is the frozen reference model, constraining deviations from the initial policy, $g:\mathbb{R}\rightarrow\mathbb{R}$ is a monotonically increasing function mapping preference scores $s_{w}$ and $s_{l}$ (for winning and losing responses $y_{w}$ and $y_{l}$) into the logit space, $\beta$ is a temperature parameter, $\mathcal{D}$ represents the dataset distribution over input-output pairs $(x,y_{w},y_{l})$.
This formulation extends the DPO framework by incorporating preference differences $g(s_{w})-g(s_{l})$, which reflect the utility of key regions identified by UV-CoT.
B.2 DPO Background and Reparameterization
DPO reformulates reward model training as a policy optimization problem by reparameterizing the reward function from Proximal Policy Optimization (PPO) [27]:
$r(x,y)=\beta\log\frac{\pi_{\theta}(y\mid x)}{\pi_{\mathrm{ref}}(y\mid x)}+\beta\log Z(x),$ (7)
where $Z(x)$ is the partition function. Substituting this into the Bradley-Terry preference model [5] yields:
$p(y_{w}\succ y_{l})$ $=\sigma\left(r(x,y_{w})-r(x,y_{l})\right)$ (8)
$=\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}\mid x)}{\pi_{\mathrm{ref}}(y_{w}\mid x)}-\beta\log\frac{\pi_{\theta}(y_{l}\mid x)}{\pi_{\mathrm{ref}}(y_{l}\mid x)}\right),$
where $\sigma$ is the sigmoid function. Maximizing the log-likelihood of this preference model leads to the naive DPO loss:
$\mathcal{L}_{\mathrm{DPO}}(\theta)$ $=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\Bigg{[}\log\sigma\Bigg{(}\beta\log\frac{\pi_{\theta}(y_{w}\mid x)}{\pi_{\mathrm{ref}}(y_{w}\mid x)}$ (9)
$\quad-\beta\log\frac{\pi_{\theta}(y_{l}\mid x)}{\pi_{\mathrm{ref}}(y_{l}\mid x)}\Bigg{)}\Bigg{]}.$
Table 6: The details of the datasets, which spans five distinct domains and includes various source datasets.
Domain Source Dataset Size Dataset Description
Text/Doc TextVQA 16k Images with text
DocVQA 35k Doc images
DUDE 15k Doc images
SROIE 4k Invoice images
Chart InfographicsVQA 15k Infographic images
General VQA Visual7W 43k Images
Flickr30k 136k Images
Relation Reasoning VSR 3k Images
GQA 88k Images
High-Resolution V Bench 238 Images
B.3 Derivation with Gumbel Distribution
To incorporate preference-weighted optimization, we define $\Delta_{r}=g(s_{w})-g(s_{l})$ and introduce Gumbel-distributed random variables $R_{w}\sim\operatorname{Gumbel}(r(x,y_{w}),1)$ and $R_{l}\sim\operatorname{Gumbel}(r(x,y_{l}),1)$. The probability that the winning response is preferred, adjusted by the preference gap, is:
$P(R_{w}-R_{l}>\Delta_{r})=\sigma\left(r(x,y_{w})-r(x,y_{l})-\Delta_{r}\right)$ (10)
$=\sigma\left(\beta\log\frac{\pi^{*}(y_{w}\mid x)}{\pi_{\mathrm{ref}}(y_{w}\mid x)}-\beta\log\frac{\pi^{*}(y_{l}\mid x)}{\pi_{\mathrm{ref}}(y_{l}\mid x)}-\Delta_{r}\right).$
This result leverages the Gumbel-max trick [20], with a similar derivation found in ODTO [1].
B.4 Proof of Gumbel-Based Preference
We first prove the foundational probability $P(R_{w}-R_{l}>0)=\sigma(\Delta_{\hat{r}_{\theta}})$, where $\Delta_{\hat{r}_{\theta}}=r_{\theta}(x,y_{w})-r_{\theta}(x,y_{l})$.
Proof: Define the random variable $I=\arg\max_{l,w}\{R_{l},R_{w}\}$. The goal is to show:
$P(I=w)=\frac{\exp(\hat{r}_{\theta}(x,y_{w}))}{\exp(\hat{r}_{\theta}(x,y_{w}))+\exp(\hat{r}_{\theta}(x,y_{l}))}.$ (11)
For notation simplicity, let $\hat{r}_{w}=\hat{r}_{\theta}(x,y_{w})$, $\hat{r}_{l}=\hat{r}_{\theta}(x,y_{l})$, and $g_{\hat{r}_{w}}\sim\operatorname{Gumbel}(\hat{r}_{w},1)$. Then:
$P(I=w)=\mathbb{E}_{m\sim g_{\hat{r}_{w}}}\left[P(R_{l}<m)\right]$ (12)
[…]
where the integral accounts for the Gumbel CDF. Let $Z=\exp(\hat{r}_{w})+\exp(\hat{r}_{l})$. The expression simplifies to:
$P(I=w)=\frac{\exp(\hat{r}_{w})}{Z}=\frac{\exp(\hat{r}_{\theta}(x,y_{w}))}{\exp(\hat{r}_{\theta}(x,y_{w}))+\exp(\hat{r}_{\theta}(x,y_{l}))},$ (14)
proving Equation (11). Extending this to the preference gap $\Delta_{r}$, we derive:
$P(R_{w}-R_{l}>\Delta_{r})$ $=1-\mathcal{F}(\Delta_{r})$ (15)
$=\frac{1}{2}-\frac{1}{2}\tanh\left(\frac{\Delta_{r}-\Delta_{\hat{r}_{\theta}}}{2}\right)$
$=\sigma(\Delta_{\hat{r}_{\theta}}-\Delta_{r}),$
where $\Delta_{\hat{r}_{\theta}}=r_{\theta}(x,y_{w})-r_{\theta}(x,y_{l})$. This completes the derivation of the UV-CoT loss in Equation (6).
Appendix C Implement Details
C.1 Datasets
To generate diverse and comprehensive image-level Chain-of-Thought (CoT) data for training Multimodal Large Language Models (MLLMs), we select nine source datasets spanning four distinct domains: Text/Doc, General Visual Question Answering (VQA), Charts, and Relation Reasoning. These domains are chosen to ensure a broad representation of visual reasoning tasks, enabling the model to develop robust CoT capabilities across varied contexts.
Before performing preference optimization, we conduct Supervised Fine-Tuning (SFT) using 10% of the labeled Visual-CoT dataset, which corresponds to approximately 25k samples, as detailed in Tab. 6. This subset is chosen to balance computational efficiency with sufficient exposure to diverse reasoning patterns, resulting in a model we denote as UV-CoT (10%). Following SFT, we perform preference optimization using a total of 249k preference data points, curated from the same nine datasets. The preference data is generated by ranking model outputs for each dataset, ensuring that the distribution across domains mirrors that of Visual-CoT [28]. Specifically, for each dataset, we maintain roughly the same proportion of preference pairs as in Visual-CoT (e.g., Text/Doc datasets contribute approximately 50% of the data, consistent with their representation in our dataset). This approach ensures that UV-CoT benefits from a balanced and comprehensive preference optimization process, enhancing its ability to prioritize key regions in visual reasoning tasks.
As shown in Tab. 6, we provide a detailed introduction to the datasets we used. For the Text/Doc domain, we include four datasets focusing on text recognition and comprehension in diverse document and image formats: DocVQA [21], TextVQA [29], DUDE [33], and SROIE [10]. These datasets provide rich text-based reasoning scenarios, such as extracting information from invoices (SROIE) or answering questions about document content (DocVQA), which are crucial for generating CoT data that involves step-by-step text interpretation.
In the General VQA domain, we select Flickr30k [25] and Visual7W [44]. These datasets are well-suited for general visual question answering, as they contain diverse images paired with questions that require understanding both visual content and textual prompts, facilitating the creation of CoT data for general reasoning tasks.
For the Charts domain, we use the InfographicsVQA dataset [22], which consists of high-resolution infographics. This dataset is particularly advantageous for training MLLMs to localize and interpret specific regions in charts, enabling the generation of CoT data that involves reasoning about data visualization elements such as legends, labels, and trends.
In the Relation Reasoning domain, we select VSR [16] and GQA [11]. These datasets are rich in spatial relational information among objects in images, making them ideal for constructing CoT data that focus on reasoning about object relationships, such as identifying relative positions or dependencies in a scene.
For high-resolution image reasoning, we use V Bench [35], which comprises 238 images from the SA-1B dataset [12] with an average resolution of 2246$\times$1582.
C.2 Evaluation
We utilize GPT-4o to assess the performance of our model due to its superior reasoning capabilities and adopt an evaluation prompt to. The prompt template is like:
Evaluating Prompt Template You are a helpful and precise assistant for checking the quality of the answer, you need to give a score to the model’s answer by referring to the standard answer, based on the given question. The full score is 1 point and the minimum score is 0 points. Please output the score in the form ‘score: <score >’. The evaluation criteria require that the closer the model’s answer is to the standard answer, the higher the score.
The meaning score, ranging from 0 to 1, reflects the semantic relevance of the model’s responses to the given prompts.
Appendix D Limitations
While our UV-CoT model demonstrates high performance across most evaluated datasets, it encounters challenges in accurately identifying anchor boxes on certain datasets, notably DocVQA [21] and InfographicsVQA [22]. These difficulties may arise due to the complex layouts, variable text densities, and noisy annotations prevalent in these datasets, which complicate the localization of relevant regions. In contrast, the ground truth (GT) boxes achieve exceptional performance on these datasets, suggesting that our model has significant untapped potential. Future research could explore advanced anchor box detection algorithms, such as incorporating adaptive thresholding or multi-scale feature extraction, to address these limitations and enhance the model’s robustness across diverse visual domains.
Appendix E Potential negative societal impacts
The potential negative societal impacts of our work align with those of other MLLMs and LLMs. While the development of UV-CoT and MLLMs advances AI capabilities, it also introduces several risks. These include heightened privacy concerns, the reinforcement of existing biases, the spread of misinformation, job displacement due to automation, and ethical challenges related to accountability, transparency, and informed consent. Addressing these issues requires responsible deployment, continuous monitoring, and the implementation of safeguards to mitigate unintended consequences.